### v1

# Diversifying the Expert Knowledge for Task-Agnostic Pruning in Sparse Mixture-of-Experts

Abstract
By increasing model parameters but activating them sparsely when performing a task, the use of Mixture-of-Experts (MoE) architecture significantly improves the performance of Large Language Models (LLMs) without increasing the inference cost. However, the memory consumption due to the growing number of experts presents a challenge to the deployment of these models in many real world settings.
Our empirical study reveals that some experts encode redundant knowledge during pre-training. We thus propose a method of grouping and pruning similar experts to improve model’s parameter efficiency. We validate the effectiveness of our method by pruning two state-of-the-art MoE models, Mixtral-8x7B and Mixtral-8x22B. Evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks. To facilitate future research, we will release our code and the pruned MoE models.
[…]
siunitx
Diversifying the Expert Knowledge for Task-Agnostic Pruning
in Sparse Mixture-of-Experts
Zeliang Zhang1111This research was conducted during the author’s internship at Microsoft Research.  Xiaodong Liu2   Hao Cheng2  Chenliang Xu1  Jianfeng Gao2 1University of Rochester 2Microsoft Research {zeliang.zhang, chenliang.xu}@rochester.edu, {xiaodl, cheng.hao, jfgao}@microsoft.com
1 Introduction
Large Language Models (LLMs) have achieved outstanding performance across various tasks by learning a large number of model parameters on large amounts of data, as shown by the scaling laws [16]. In addition to increasing the depth of neural network models, widening neural networks by using the sparsely-activated mixture-of-experts (MoE) architecture is also proved effective. MoE widens the feed-forward network (FFN) layer (one expert) by having multiple parallel FFNs (experts). During forward propagation, only a subset of these experts is activated. Thus, compared to dense models, MoE models achieve better end-task performance and generalize better to new tasks without increasing computation costs. Notable examples of MoE models include Switch Transformer [9], Mixtral-MoE [14], and Uni-MoE [22].
Despite significant progress in developing wider and deeper MoE LLMs, the increased memory consumption due to larger model sizes (i.e., increased number of experts) poses a substantial challenge to the deployment of these models in real-world settings. For example, storing and loading Mixtral-8x7B, which has 8 experts in each of its 32 layers, requires approximately $88$ GB. The MoE layers constitute the majority of the parameters. Adding or removing even one expert in each layer can significantly impact overall memory cost and model performance. For example,  [27] shows that randomly dropping $2$ experts in each MoE layer reduces the memory cost by $21$ GB, and decreases model performance by $7\%$ on the MMLU benchmark [12]. In this study, we strive to seek the best trade-off between memory efficiency and task performance by identifying an optimal set of experts in each MoE layer to prune.
There are many recent studies on pruning MoE models. One line of work utilizes task-specific information to prune irrelevant experts. For example, Chen et al. [3] prune the less frequently visited experts based on experiments on a range of tasks. Chowdhury et al. [5] find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage. Li et al. [21] merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning. Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks. He et al. [11] explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop. In comparison, Lu et al. [27] enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance. We illustrate the difference among these works in fig. 1. Although pruning MoE models in task-agnostic settings is of great practical value, this area has not been fully explored.
Figure 1: Removing several experts from the original MoE layer would not cause model collapse but improve efficiency. Prior works often utilize expert access information to filter out unimportant experts. In our work, we first group different experts with similar knowledge in the feature space, then merge them along with the routers to prune the MoE layer. This post-processing approach allows us to diversify the features of each MoE layer, thereby preserving the knowledge of the original large models as much as possible while reducing computation and storage consumption.
In this work, we explore how to prune MoE models in a task-agnostic fashion. Our study is motivated by the finding that, given the same input, many experts respond similarly, indicating that these experts likely encode similar knowledge, and thus are somewhat redundant. We propose a method to improve model parameter efficiency by pruning redundant experts in two stages. As shown in fig. 1, we first identify and group similar experts in the feature space. Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers. We validate the effectiveness of our method by pruning experts for two state of the art MoE models, Mixtral-8x7B and Mixtral-8x22B. Evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
2 Related Work
Sparse MoEs. The Mixture-of-Experts (MoE) structure is firstly applied in classical machine learning models by Jacobs et al. [13] and Jordan and Jacobs [15], then widely used in various deep learning models [38, 30, 39]. Recently, some works employ the MoE to scale the capacity of transformer-based models, especially the large language models [33, 19, 42]. It adapts the original large feed-forward network (FFN) in each transformer block into multiple smaller FFNs, forming an expert layer with a router that computes the weighted output of each MoE layer. Sparse MoEs were first proposed by Fedus et al. [9]. In this approach, only a few experts are activated in each layer, accelerating training and inference while significantly increasing the number of parameters for greater model capacity. Many sparse MoEs have been developed and open-sourced within the AI community, such as Switch Transformer [9], Mixtral-8B [14], and Uni-MoE [22]. Recent studies also indicate that neural networks with the MoE structure exhibit better generalization ability compared to dense models [40, 20].
Model Pruning. Model pruning involves removing unimportant parameters from a well-trained neural network to balance task performance and computational efficiency [25, 36, 24]. Pruning techniques can be categorized into unstructured pruning [23, 35, 29], which introduces sparsity in the weight matrix by setting some parameters to zero, and structured pruning [18, 8, 34], which removes entire neurons, layers, or blocks, reducing redundancy and being more suitable for acceleration on GPUs [4]. Many efforts have been made to leverage model pruning techniques to reduce the memory consumption of neural networks, spanning a range of models from conventional architectures like CNNs [28], RNNs [41], and LSTMs [7] to modern large models such as Llama [37] and Stable-Diffusion [2].
While the large amount of parameters in sparse MoEs benefits the model’s capacity to achieve good performance at the pre-training stage, the increasing memory consumption causes great challenges to fine-tuning different downstream tasks. In this paper, we work on pruning the sparse MoEs to reduce redundant experts in the task-agnostic setting, which enhances the computational and memory efficiency throughout the fine-tuning process, and scalability of deploying these models.
3 Methodology
Notations. We denote $F(\cdot;\Theta,W,K)$ as an MoE layer, where $\Theta=\{\theta_{1},\theta_{2},...,\theta_{N}\}$ indicates the parameter set for $N$ experts $\{f_{n}(\cdot;\theta_{n})\}_{n=1}^{N}$, $W\in\mathbb{R}^{N\times d}$ indicates the routing weights for each expert and $K$ indicates the number of experts selected for each token.
For a new token $x\in\mathbb{R}^{d}$, we first compute the routing logits to score the match between $x$ and the $n$-th expert:
$p_{n}(x)=\frac{e^{W_{n}x}}{\sum_{t=1}^{N}e^{W_{t}x}},$ (1)
where $W_{t}\in\mathbb{R}^{d}$ is the routing policy of $t$-th expert. Then, we take the experts $\{f_{n}(\cdot;\theta_{n})\}_{n=i_{1}}^{i_{K}}$ with top-$K$ logits $\{l_{n}\}_{n=i_{1}}^{i_{K}}$ as the selected target experts. Last, we compute the weighted average of the outputs from selected experts as the final output of the MoE layer,
$y=\sum\limits_{n=i_{1}}^{i_{K}}p_{n}(x)\cdot f_{n}(x;\theta_{n}).$ (2)
3.1 Task Definition
Several prior works [27, 11, 21] have identified that removing some experts in different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks. The remaining question is which expert(s) should be pruned in the MoE layer.
Specifically, for a MoE LLM $\mathcal{M}(\cdot;\mathcal{F})$, which consists of multiple MoE layers =$\mathcal{F}=\{F^{l}(\cdot;\Theta^{l},W^{l},K)\}_{l=1}^{L}$, we want to prune $r$ experts from the total $N$. This leads to the expert modules $\Theta$ being reduced to $\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{r}}\}$, and the routing matrix $W$ being reduced to $\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{r}}\}$, where $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{r}\}$ is the set of indices to delete. We aim to find the pruned $\hat{\mathcal{M}}$ by deleting the optimal indices set $s^{l}$ in different MoE layers $l$ that satisfy the following optimization objective:
$\min\limits_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{\mathcal{M}}(x;\hat{\mathcal{F}}),y),$ (3)
where $\mathcal{L}$ is the loss function, $\hat{\mathcal{M}}(\cdot;\hat{\mathcal{F}})$ is the pruned LLM by deleting $r$ experts from each MoE layer $F$ in $\mathcal{F}$, and $\mathcal{D}$ is the dataset.
This task is particularly challenging as this is an NP-hard search problem with $(C^{r}_{N})^{L}$ possible solutions. In this work, we propose to estimate expert similarity in the feature space for greedy structured pruning within each MoE layer to address this problem.
Figure 2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under two kernel-based CKA criteria (Linear and RBF). A darker color indicates a greater similarity between experts.
3.2 Expert Similarity in MoE layer
Motivation: Intuitively, the more independent the experts are within the same MoE layer, the more diverse the representations learned by each expert. This diversity makes it harder to replace or drop any of the experts. Conversely, the redundancy of shared knowledge learned by different experts provides opportunities for model pruning.
We employ the Centered Kernel Alignment (CKA) [17] as the criteria to evaluate the similarity between experts in each MoE layer. Specifically, given a batch of inputs $\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ between two experts $f_{i}$ and $f_{j}$ is computed as follows,
$\rho_{ij}=\frac{\text{HSIC}(K^{i},K^{j})}{\text{HSIC}(K^{i},K^{i})\cdot\text{HSIC}(K^{j},K^{j})},$ (4)
where $\text{HSIC}(K^{i},K^{j})=\frac{1}{(s-1)^{2}}\text{tr}(K^{i}HK^{j}H)$, $K^{i}_{mn}=k(f_{i}(x_{m}),f_{i}(x_{n}))$, $H=I-\frac{1}{s}11^{T}$, and $k(\cdot,\cdot)$ is the kernel function.
Here, we compute the expert similarity of Mixtral-8Bx7B using randomly selected $32$ samples from the pre-trained dataset C4 [32]. We respectively use the linear kernel and RBF kernel in similarity computation of eq. 4, where the use of linear kernel actually corresponds to the dot product similarity between the expert outputs.
As illustrated in Fig. 2, the darker the cell, the more similar the two related experts. The two kernel-based distance measures demonstrate similar changes in expert similarity across different layers. There are varying moderate ratios of similar paired experts in different MoE layers, suggesting the plausibility of pruning the MoE layer with many experts down to fewer experts. For example, in the first layer, both the linear kernel and the RBF kernel indicate that the $2$-nd to $5$-th experts exhibit great similarity (greater than 0.7) with each other, suggesting that one of them can be pruned without significantly degrading model performance.
Remark.
We notice that a concurrent work [26] also studies expert similarity, but focuses on the cosine similarity between expert weights. Our work examines the differences between experts using a more generalized metric based on the correlation between inputs and outputs, and obtains a different observation: even without considering the last layer, expert similarity does not gradually decrease with increasing depth of layers. For example, we observe higher pairwise expert similarity at the $11$-th layer compared to the $6$-th layer, which is cross-verified by both distance metrics. Additionally, the experts in the $26$-th layer also demonstrate a higher similarity with others compared with shallow layers. We provide the full evaluation results on visualizing the expert similarity of MoE layers in Mixtral-8x7B and Mixtral-8x22B in appendix A.
3.3 Discover and Merge Similar Experts on Feature Space
Figure 3: Overview of our method. We first leverage different metrics to obtain the expert representation, then compute the pairwise distance to get the expert similarity matrix. Based on the expert similarity matrix, we can group the experts with shared knowledge in the same cluster, which can be merged on the weight space for MoE pruning.
Having identified the existence of moderate expert similarity in each MoE layer, we propose pruning the experts within each MoE layer by merging experts with shared knowledge. We present an overview of our method in fig. 3. Suppose we want to reduce the number of experts from $N$ to $r$, the details of our approach are outlined as follows.
Discovering Similar Experts from the Expert Graph. We construct a unidirectional graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ to group similar experts into the same clusters, where we denote $\mathcal{V}$ as the set of nodes and $\mathcal{E}$ as the connections between nodes.
First, for each expert $f_{i}$, we compute its feature representation $\mathcal{M}(f_{i})$, where $\mathcal{M}(\cdot)$ is a specialized function to represent $f_{i}$.
Second, we compute the pairwise distance based on expert representation and obtain the disparity matrix $D\in\mathbb{R}^{N\times N}$, where $D_{ij}=\text{DistanceFunction}(\mathcal{M}(f_{i}),\mathcal{M}(f_{j}))$.
Then, we initialize the $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the distance matrix $D$.
Last, we leverage the graph partition algorithm to split $\mathcal{G}$ into $r$ subgraphs $\{\mathcal{G}_{i}\}_{i=1}^{r}$. The experts indexed by $\mathcal{V}_{i}$ of $\mathcal{G}_{i}$ shares the most similar knowledge with each other, and shows much difference with experts indexed by $\mathcal{V}_{t}$ of $\mathcal{G}_{t}$, $t\neq i$. To achieve this goal, we adopt a two-step strategy. At the first step, we partition the $\mathcal{V}$ of $\mathcal{G}$ into two subsets $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, where $\mathcal{V}_{1}$ has $r-1$ nodes and $\mathcal{V}_{2}$ has $G-r+1$ nodes. Then, we split $\mathcal{V}_{1}$ into $r-1$ subsets, $\mathcal{V}_{1,1},...,\mathcal{V}_{1,r-1}$. The splitting process can be formulated as follows,
$\min\sum_{i,j\in\mathcal{V}}x_{i}x_{j}D_{ij}-\sum_{i,j\in\mathcal{V}}x_{i}(1-x_{j})D_{ij}$ (5)
$-\sum_{i,j\in\mathcal{V}}(1-x_{i})(1-x_{j})D_{ij},$
[…]
where $x_{i}=1$ if node $i$ in group $\mathcal{V}_{1}$, and $x_{i}=0$ otherwise. The first term is used to minimize the distance between nodes in $\mathcal{V}_{1}$, the second term is usd to maximize the distance of nodes between $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, and the third term is ued to maximize the distance of nodes in $\mathcal{V}_{2}$.
When pruning $G$ experts to $r$ experts, there are $r-1$ experts keeping unaffected in the entire process. Our graph partition strategy can preserve the most knowledge while diversifying the expert features in each MoE layers.
Merging Similar Experts with Routers. To preserve the different knowledge learned by different models clustered in the same group, we merge these experts as well as their routers on the weight space as follows,
$\hat{\theta}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}\theta_{\mathcal{V}_{n}(i)},$ (6)
$\hat{W}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$
where $\mathcal{V}_{n}$ is the set of similar experts in the $n$-th cluster $\mathcal{G}_{n}$, and we have $\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}=1$. We update the MoE layer $F$ by respectively replacing all the experts grouped in $\mathcal{G}_{n}$ by a single FFN expert layer $f(\cdot;\hat{\theta}_{n})$ and corresponding routing weights $\hat{W}_{n}$, $n=1,2,...,N-r$.
3.4 Practical Implementation
While we have developed a general strategy to prune sparse MoEs, several practical questions arise during implementation:
Q1: How to decide $\mathcal{M}(\cdot)$ and perform task-agnostic pruning on sparse MoEs?
Intuitively, given the same input, similar experts should generate similar outputs. Thus, we propose to use the mean squared error on the outputs of two experts to evaluate the similarity between them.
We follow previous work [27] to select few samples from pre-trained dataset C4 [32] as the calibration dataset to compute the expert feature representation at the third step for similar experts discovery. The C4 dataset comprises multiple domains of contents, which can be a good surrogate to simulate the performance of the task-agnostic pruning.
However, with the limited size of the calibration dataset, the estimation of expert similarity based on the aforementioned expert representation could be inaccurate due to biased sampling. To improve the generalization ability at the discovery stage, we randomly mix up the token embedding for data augmentation and then use the augmented data for computing the expert representation and similarity. In practice, we iteratively update the expert similarity matrix by continuously shuffling and mixing up the data for full augmentation.
Q2: Is it possible to prune sparse MoEs without using any data?
Experts in the MoE layer of current models typically have the same structure; for instance, in Mixtral-8x7B, each expert is a two-layer FFN with SiLU activation. By training on different routed tokens, the experts usually learn different distributions during the training process. This results in a disparity between the learned feature spaces of different experts, which is also potentially reflected by their weights.
Based on this analysis, we can use the weights as the expert representation to discover similar experts for pruning. In our implementation, we consider two strategies: (1) directly using the vectorized weights, which does not account for the structure of experts, and (2) leveraging the local linearity of neural networks to compute the approximated weight matrix, namely the surrogate weights in our paper.
Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$. For input $x$, the output is computed as $f(x;\theta)=\theta_{2}(\sigma(\theta_{1}x)\cdot\theta_{3}x)$. Then, the vectorized weight is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$.
Q3: How do we determine the value of $\alpha$ for merging similar experts?
Two strategies are explored in our implementation to merge similar experts.
One strategy is to simply merge experts and gates with the same coefficients, i.e., $\alpha_{1}=\alpha_{2}=...=\alpha_{|\mathcal{V}_{n}|}=\frac{1}{|\mathcal{V}_{n}|}$, which encourages experts contributing equally.
The other strategy is to learn the merging coefficients for different experts towards minimizing the following loss function,
($\mathcal{L}(\{\alpha_{n}\}_{n=1}^{|\mathcal{V}_{n}|})=\|y-F(x;\hat{\Theta},\hat{W},K)\|,$ (7)
[…]
$\hat{W}_{n}=\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$
where the ground-truth is the output of original MoE layer $y=F(x;\Theta,W,K)$, and we jointly optimize $\lambda$ and $\alpha$ for different merging groups.
4 Evaluation
Table 1: Results on pruning the Mixtral-8x7B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Augmented: prunable experts discover using augmented data and merging with uniform coefficients; 3) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq. 7. The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
Router-guided [21] 51.8/24.8 60.5/26.5 46.9/24.7 60.5/25.0 82.6/39.9 32.0/11.6 70.4/50.9 57.8/29.1
Count-guided [11] 49.2/36.9 59.7/45.6 45.0/35.1 58.2/43.4 77.2/76.6 33.0/26.4 56.6/55.9 54.1/45.7
Enumerate [27] 52.4/43.5 66.4/52.7 49.0/40.4 63.7/43.5 84.0/80.8 32.6/28.8 71.1/66.4 59.9/50.8
Ours Vectorized $\bf{\theta}$ 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
Surrogate $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
Augmented 56.2/47.8 72.7/57.2 52.1/42.7 68.1/54.0 85.6/83.2 32.8/28.6 68.6/66.5 62.3/54.3
Learn 56.0/48.0 73.1/57.0 52.4/43.3 68.2/54.6 86.4/83.3 31.4/28.5 69.3/67.1 62.4/54.5
4.1 Experiment Setup
Studied Models. We take pruning Mixtral-8x7B and Mixtral-8x22B [14] as the example. The Mixtral-8x7B has $32$ sparse MoE-involved layers, in each there are $8$ experts. The hidden dimension for token embedding is $4,096$, and the intermediate feature dimension in each FFN of MoEs is $14,336$. The Mixtral-8x22B is similar with Mixtral-8x7B, but with $56$ sparse MoE layers. It has the token embedding dimension of $6,144$ and the intermediate feature dimension of $16,384$. During the inference, each token will select $2$ experts in each MoE layer.
Compared Baselines. We take three advanced MoE pruning methods [11, 21, 27] as our baseline for comparison. Among them, router-guided merging [21] is initially designed for task-specific MoE pruning, where we set the target dataset as the samples from the pre-training dataset. For task-agnostic methods, we select the frequency-based pruning method Expert Trimming [27], also the count-guided strategy and loss-based pruning method [27], namely the Enumerate in our paper. Under the task-agnostic pruning setting, we disable all the fine-tuning stage of all methods for fair comparison.
Our Approaches. We use four metrics to prune the model, including 1) Vectorized $\theta$: concact the vectorized model parameter $\theta$ together as the expert representation for grouping, and apply the uniform merging on the experts and routers. 2) Surrogate $\theta$: multiply the model parameters together as the weight for one-layer surrogate FFN and use it as the expert representation. The merging strategy is uniform merging. 3) Augmented: compute the expert representation based on the expert outputs on the augmented inputs. The merging strategy is uniform merging. 4) Learn: Use the same grouping strategy with Augmented but learn to merge for experts and routers.
Evaluation Datasets. The open-sourced Language Model Evaluation Harness library [10] is used to evaluate the performance. We select four zero-shot tasks, including MMLU [12], BoolQ [6], OpenBookQA [31], and RTE [1]. Among these tasks, MMLU consists of $57$ subtasks, where we present four groups, namely the humanities, social science, stem and other.
Implementations. We randomly select $128$ samples from the pre-training dataset C4 [32] as the calibration dataset. For the prunable candidate group discovery, we use the mean square error to evaluate the similarity between experts. For the learnable merging, we initialize the merging coefficients $\alpha$ as the uniform and set $\lambda$ as $1$. In our experiments, we utilize four NVIDIA A100 GPUs, each with 80GB of memory.
4.2 Pruning the Mixtral Architecture
We present the results of pruning the Mixtral architecture, including Mixtral-8x7B and Mixtral-8x22B, from 8 experts to 6 and 4 experts in each layer, respectively. The results on the two models are depicted in table 1 and table 2 respectively.
Results on Mixtral-8x7B. We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average. Compared with count-guided strategy [11] which just drops the experts less visited, router-guided strategy [21] has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge. Besides, compared with count-guided and enumerate strategies [27] which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers.
From the results of our strategies, we can also notice two facts: 1) Both Vectorized $\theta$ and Surrogate $\theta$ strategies surpass all the other data-involved pruning baseline methods. This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning. 2) Compared with only using $\theta$, the involvement of $data$ in the pruning process can help better candidates for merging and pruning for improving the performance. Besides, though the learning strategy can bring a further improvement in the average performance, the computation cost is larger compared with using uniform coefficients for merging.
Results on Mixtral-8x22B. With more parameters introduced in Mixtral-8x22B, the search for optimal experts to drop and merge becomes more challenging. Unlike the results on pruning Mixtral-8x7B, the strategy using the surrogate $\theta$ achieves state-of-the-art results, with only a minor performance drop of $2.8\%$ on average compared to the full model.
At the same time, it should be noted that the use of data doesn’t help boost the pruning performance as well, where the augmented and learning-based method has an average performance drop of $1.05\%$ and $1.0\%$, respectively. Also, the routing similarity-based pruning suffers model collapse under both two settings. We argue it is due to the heavy imbalance between the used data and parameter space, leading to the computation of expert similarity over-fitted on the small calibration dataset.
Table 2: Results on pruning the Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$, 2) Augmented, and 3) Learn. The first column indicate the results on pruned model with $6$ experts, while the second column present the result of $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
Router-guided [21] 27.3/22.7 25.4/25.8 24.4/24.0 27.9/23.4 62.8/62.7 12.8/13.0 54.2/49.5 33.5/31.6
Count-guided [11] 58.0/45.7 74.9/57.7 54.1/42.0 70.2/45.7 81.5/74.4 35.2/27.0 69.3/57.4 63.3/50.0
Enumerate [27] 60.4/53.9 78.0/67.2 59.5/52.3 73.0/64.2 87.4/80.5 35.0/31.1 70.1/67.9 66.2/59.6
Ours Vectorized $\bf{\theta}$ 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
Surrogate $\bf{\theta}$ 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
Augmented 62.3/57.8 78.5/69.7 60.2/51.3 73.4/64.2 87.6/83.1 35.8/33.2 71.1/68.1 67.0/61.1
Learn 61.4/57.9 78.3/70.1 61.2/51.4 72.8/65.0 88.2/84.9 35.6/32.7 70.5/67.3 66.9/61.3
4.3 Empirical Analysis on Expert Hints
While expert pruning brings a significant speedup by reducing the parameters, what changes are introduced to different MoE layers? To answer this question, we analyze the change of expert hints on the calibration dataset C4.
Figure 4: Statistics of the visiting frequency for all experts in different MoE layers.
The statistic results are shown in fig. 4. We can see that in most MoE layers in Mixtral-8x7B, many pairs of experts have similar hints. In contrast, pruning differentiates the hints of experts. Compared to directly using hints as the pruning goal, using expert knowledge as the pruning criterion results in more significant changes in hints. Additionally, the merging operation on the gate aggregates the new expert (last column) more tokens, further increasing the hint differences.
4.4 Ablation Study and Discussion
On the number of samples. In our work, we augment the data to improve the accuracy of similarity estimation. While this technique is typically used in scenarios with limited data, a natural question arises:
Would increasing the amount of data help preserve performance after pruning?
Table 3: Evaluation results of different pruning setting with ranging the number of samples for expert similarity estimation. The augmentation is disabled in this experiment.
# of samples MMLU BoolQ OpenBookQA RTE Avg.
128 0.615 0.856 0.328 0.686 0.621
256 0.614 0.854 0.332 0.697 0.624
512 0.617 0.854 0.334 0.704 0.627
2,048 0.612 0.856 0.322 0.675 0.616
We conduct experiments with different numbers of samples to answer this question. As shown in table 3, when we increase the number of samples from $128$ to $512$, we can see a noticeable performance improvement with up to $0.6\%$. It should be noted that when we scale up the number of samples to $2,048$, the average performance dramatically drops to $61.6\%$.
We attribute this phenomenon to the use of mean operation in our expert similarity estimation. On the one hand, the mean operation keeps the feature dimension for efficient pairwise distance computation. On the other hand, it naturally downgrades the precision of function similarity computation by “round-off” error compared with other per sample-based distance metrics like CKA used in section 3.2.
On the choice of distance function. We also considered other distance functions, such as cosine similarity, to evaluate expert similarity during the discovery stage. However, it performed worse compared to our mean-squared-error function, showing a drop of $0.9\%$ on MMLU, $3.3\%$ on RTE, and $0.2\%$ on BoolQ.
On merging the routing policy. One relevant work to ours is [21], which identifies and merges similar experts based on routing policy, setting the routing policy of the merged expert to that of the most frequently visited expert. While we identify similar experts based on expert knowledge, such as weights and features, we merge not only the experts but also the routing policies. To demonstrate the importance of merging routers, we evaluate the performance of our pruning method without merging the routing policies. As shown in table 4, this results in performance degradation across all tasks and different numbers of experts, highlighting the crucial role of simultaneously merging the routers and experts.
Table 4: Evaluation results of our augmentation-based pruning method without merging the routers.
# of experts MMLU BoolQ OpenBookQA RTE Avg.
4 0.491 0.833 0.274 0.664 0.565
6 0.604 0.847 0.312 0.675 0.609
5 Conclusion and Future Work
In this paper, we address the practical problem of reducing the number of experts in MoE layers to accelerate the inference and fine-tuning process. We propose discovering similar experts at the feature level and then merging them in the weight space for MoE pruning while preserving as much original expert knowledge as possible. This approach allows the MoE layer to maintain diverse experts with different knowledge, thereby efficiently reducing redundancy.
However, current methods do not account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers. Our analysis reveals that some experts have low visitation frequencies, indicating significant potential for further pruning. Future work will aim to push the compression of MoE layers to the limit.
6 Acknowledgments
We would like to thank Lifeng Li from Microsoft for setting up GPU machines.
Appendix A Evaluation on the expert similarity
Following the same setting in section 3.2, we conduct experiments on evaluating the expert similarity in all MoE layers of Mixtral-8x7B and Mixtral-8x22B. The results are respectively depicted in fig. A1 and table 2. We summarize the observation as follows,
  • Most MoE layers in the two Mixtral models contain significant expert redundancy.
  • The most redundant MoE layers are located in the first and last several layers, while the experts in the intermediate MoE layers learn more diverse features.
Figure A1: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under the linear kernel-based CKA criteria.
Figure A2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x22B under the linear kernel-based CKA criteria.