### v2

# Diversifying the Expert Knowledge for Task-Agnostic Pruning in Sparse Mixture-of-Experts

Abstract
By increasing model parameters but activating them sparsely when performing a task, the use of Mixture-of-Experts (MoE) architecture significantly improves the performance of Large Language Models (LLMs) without increasing the inference cost. However, the memory consumption due to the growing number of experts presents a challenge to the deployment of these models in many real-world settings.
Our empirical study reveals that some experts encode redundant knowledge during pre-training. We thus propose a method of grouping and pruning similar experts to improve the model’s parameter efficiency. We validate the effectiveness of our method by pruning three state-of-the-art MoE architectures, including Mixtral, Deepseek-MoE, and Qwen. The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks. We will release our code to facilitate future research.
[…]
siunitx
Diversifying the Expert Knowledge for Task-Agnostic Pruning
in Sparse Mixture-of-Experts
Zeliang Zhang1111This research was conducted during the author’s internship at Microsoft Research.  Xiaodong Liu2   Hao Cheng2  Chenliang Xu1  Jianfeng Gao2 1University of Rochester 2Microsoft Research {zeliang.zhang, chenliang.xu}@rochester.edu, {xiaodl, cheng.hao, jfgao}@microsoft.com
1 Introduction
Large Language Models (LLMs) have achieved outstanding performance across various tasks by learning a large number of model parameters on large amounts of data, as shown by the scaling laws [16]. In addition to increasing the depth of neural network models, widening neural networks by using the sparsely-activated mixture-of-experts (MoE) architecture is also proved effective. MoE widens the feed-forward network (FFN) layer (one expert) by having multiple parallel FFNs (experts). During forward propagation, only a subset of these experts is activated. Thus, compared to dense models, MoE models achieve better end-task performance and generalize better to new tasks without increasing computation costs. Notable examples of MoE models include Switch Transformer [9], Mixtral-MoE [14], and Uni-MoE [22].
Despite significant progress in developing wider and deeper MoE LLMs, the increased memory consumption due to larger model sizes (i.e., increased number of experts) poses a substantial challenge to the deployment of these models in real-world settings. For example, storing and loading Mixtral-8x7B, which has 8 experts in each of its 32 layers, requires approximately $88$ GB. The MoE layers constitute the majority of the parameters. Adding or removing even one expert in each layer can significantly impact overall memory cost and model performance. For example,  [26] shows that randomly dropping $2$ experts in each MoE layer reduces the memory cost by $21$ GB, and decreases model performance by $7\%$ on the MMLU benchmark [12]. In this study, we strive to seek the best trade-off between memory efficiency and task performance by identifying an optimal set of experts in each MoE layer to prune.
There have been several studies on pruning MoE models. One line of work utilizes task-specific information to prune irrelevant experts. For example, Chen et al. [3] prune the less frequently visited experts based on experiments on a range of tasks. Chowdhury et al. [5] find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage. Li et al. [21] merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning. Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks. He et al. [11] explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop. In comparison, Lu et al. [26] enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance. We illustrate the difference among these works in fig. 1. Although pruning MoE models in task-agnostic settings is of great practical value, this area has not been fully explored.
Figure 1: Removing several experts from the original MoE layer would not cause model collapse but improve efficiency. Prior works often utilize expert access information to filter out unimportant experts. In our work, we first group different experts with similar knowledge in the feature space, then merge them along with the routers to prune the MoE layer. This post-processing approach allows us to diversify the features of each MoE layer, thereby preserving the knowledge of the original large models as much as possible while reducing computation and storage consumption.
In this work, we explore how to prune MoE models in a task-agnostic fashion. Our study is motivated by the finding that, given the same input, many experts respond similarly, indicating that these experts likely encode similar knowledge, and thus are somewhat redundant. We propose a method to improve model parameter efficiency by pruning redundant experts in two stages. As shown in fig. 1, we first identify and group similar experts in the feature space. Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers. We validate the effectiveness of our method by pruning experts for three state-of-the-art MoE architectures, including Mixtral (Mixtral-8x7B and Mixtral-8x22B), Deepseek-MoE (Deepseek-MoE-16B) and Qwen (Qwen2-57B-14A). The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
2 Related Work
Sparse MoEs. The Mixture-of-Experts (MoE) structure is firstly applied in classical machine learning models by Jacobs et al. [13] and Jordan and Jacobs [15], then widely used in various deep learning models [39, 29, 40]. Recently, some works employ the MoE to scale the capacity of transformer-based models, especially the large language models [32, 19, 43]. It adapts the original large feed-forward network (FFN) in each transformer block into multiple smaller FFNs, forming an expert layer with a router that computes the weighted output of each MoE layer. Sparse MoEs were first proposed by Fedus et al. [9]. In this approach, only a few experts are activated in each layer, accelerating training and inference while significantly increasing the number of parameters for greater model capacity. Many sparse MoEs have been developed and open-sourced within the AI community, such as Switch Transformer [9], Mixtral-8B [14], and Uni-MoE [22]. Recent studies also indicate that neural networks with the MoE structure exhibit better generalization ability compared to dense models [41, 20].
Model Pruning. Model pruning involves removing unimportant parameters from a well-trained neural network to balance task performance and computational efficiency [25, 36, 24]. Pruning techniques can be categorized into unstructured pruning [23, 34, 28], which introduces sparsity in the weight matrix by setting some parameters to zero, and structured pruning [18, 8, 33], which removes entire neurons, layers, or blocks, reducing redundancy and being more suitable for acceleration on GPUs [4]. Many efforts have been made to leverage model pruning techniques to reduce the memory consumption of neural networks, spanning a range of models from conventional architectures like CNNs [27], RNNs [42], and LSTMs [7] to modern large models such as Llama [38] and Stable-Diffusion [2].
While the large amount of parameters in sparse MoEs benefits the model’s capacity to achieve good performance at the pre-training stage, the increasing memory consumption causes great challenges to fine-tuning different downstream tasks. In this paper, we work on pruning the sparse MoEs to reduce redundant experts in the task-agnostic setting, which enhances the computational and memory efficiency throughout the fine-tuning process, and scalability of deploying these models.
3 Methodology
Notations. We denote $F(\cdot;\Theta,W,K)$ as an MoE layer, where $\Theta=\{\theta_{1},\theta_{2},...,\theta_{N}\}$ indicates the parameter set for $N$ experts $\{f_{n}(\cdot;\theta_{n})\}_{n=1}^{N}$, $W\in\mathbb{R}^{N\times d}$ indicates the routing weights for each expert and $K$ indicates the number of experts selected for each token.
For a new token $x\in\mathbb{R}^{d}$, we first compute the routing logits to score the match between $x$ and the $n$-th expert:
$p_{n}(x)=\frac{e^{W_{n}x}}{\sum_{t=1}^{N}e^{W_{t}x}},$ (1)
where $W_{t}\in\mathbb{R}^{d}$ is the routing policy of $t$-th expert. Then, we take the experts $\{f_{n}(\cdot;\theta_{n})\}_{n=i_{1}}^{i_{K}}$ with top-$K$ logits $\{l_{n}\}_{n=i_{1}}^{i_{K}}$ as the selected target experts. Last, we compute the weighted average of the outputs from selected experts as the final output of the MoE layer,
$y=\sum\limits_{n=i_{1}}^{i_{K}}p_{n}(x)\cdot f_{n}(x;\theta_{n}).$ (2)
3.1 Task Definition
Previous works [26, 11, 21] have identified that removing some experts from different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks. The remaining question is which expert(s) should be pruned in the MoE layer.
Specifically, for the $l$-th MoE layer $F^{l}(\cdot;\Theta^{l},W^{l},K)$ in the given LLM $\mathcal{M}(\cdot;\mathcal{F})$, we want to prune $r$ experts from the total of $N$, $l=1,2,...,L$. This leads to the expert modules $\Theta$ being reduced to $\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{r}}\}$, and the routing matrix $W$ being reduced to $\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{r}}\}$, where $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{r}\}$ is the set of indices to delete. We aim to find the pruned $\hat{\mathcal{M}}$ by deleting the optimal indices set $s^{l}$ in different MoE layers $l$ towards the following optimization objective:
$\min\limits_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{\mathcal{M}}(x;\hat{\mathcal{F}}),y),$ (3)
where $\mathcal{L}$ is the loss function, $\hat{\mathcal{M}}(\cdot;\hat{\mathcal{F}})$ is the new LLM with pruned MoE layer $\hat{\mathcal{F}}$, and $\mathcal{D}$ is any dataset under the task-agnostic setting.
This task is particularly challenging, as this is an NP-hard search problem with $(C^{r}_{N})^{L}$ possible solutions. In this work, we propose to estimate the expert similarity in the feature space for greedy structured pruning within each MoE layer to address this problem.
Figure 2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under two kernel-based CKA criteria (Linear and RBF). A darker color indicates a greater similarity between experts.
3.2 Expert Similarity in MoE layer
Motivation: Intuitively, the more independent the experts are within the same MoE layer, the more diverse the representations learned by each expert. This diversity makes it harder to replace or drop any of the experts. Conversely, the redundancy of shared knowledge learned by different experts provides opportunities for model pruning.
We employ the Centered Kernel Alignment (CKA) [17] as the criteria to evaluate the similarity between experts in each MoE layer. Specifically, given a batch of inputs $\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ between two experts $f_{i}$ and $f_{j}$ is computed as follows,
$\rho_{ij}=\frac{\text{HSIC}(K^{i},K^{j})}{\text{HSIC}(K^{i},K^{i})\cdot\text{HSIC}(K^{j},K^{j})},$ (4)
where $\text{HSIC}(K^{i},K^{j})=\frac{1}{(s-1)^{2}}\text{tr}(K^{i}HK^{j}H)$, $K^{i}_{mn}=k(f_{i}(x_{m}),f_{i}(x_{n}))$, $H=I-\frac{1}{s}11^{T}$, and $k(\cdot,\cdot)$ is the kernel function.
Here, we compute the expert similarity of Mixtral-8Bx7B using randomly selected $32$ samples from the pre-trained dataset C4 [31]. We respectively use the linear kernel and RBF kernel in similarity computation of eq. 4, where the use of linear kernel actually corresponds to the dot product similarity between the expert outputs.
As illustrated in Fig. 2, the darker the cell, the more similar the two related experts. The two kernel-based distance measures demonstrate similar changes in expert similarity across different layers. There are varying moderate ratios of similar paired experts in different MoE layers, suggesting the plausibility of pruning the MoE layer with many experts down to fewer experts. For example, in the first layer, both the linear kernel and the RBF kernel indicate that the $2$-nd to $5$-th experts exhibit great similarity (greater than 0.7) with each other, suggesting that one of them can be pruned without significantly degrading model performance.
3.3 Discover and Merge Similar Experts on Feature Space
Figure 3: Overview of our method. We first leverage different metrics to obtain the expert representation, then compute the pairwise distance to get the expert similarity matrix. Based on the expert similarity matrix, we can group the experts with shared knowledge in the same cluster, which can be merged on the weight space for MoE pruning.
Having identified the existence of moderate expert similarity in each MoE layer, we propose pruning the experts within each MoE layer by merging experts with shared knowledge. We present an overview of our method in fig. 3. Suppose we want to reduce the number of experts from $N$ to $r$, the details of our approach are outlined as follows.
Discovering Similar Experts from the Expert Graph. We construct a unidirectional graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ to group similar experts into the same clusters, where we denote $\mathcal{V}$ as the set of nodes and $\mathcal{E}$ as the connections between nodes.
First, for each expert $f_{i}$, we compute its feature representation $\mathcal{M}(f_{i})$, where $\mathcal{M}(\cdot)$ is a specialized function to represent $f_{i}$.
Second, we compute the pairwise distance based on expert representation and obtain the disparity matrix $D\in\mathbb{R}^{N\times N}$, where $D_{ij}=\text{DistanceFunction}(\mathcal{M}(f_{i}),\mathcal{M}(f_{j}))$.
Then, we initialize $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the distance matrix $D$.
Last, we split $\mathcal{G}$ into $r$ subgraphs $\{\mathcal{G}_{i}\}_{i=1}^{r}$ to group similar experts. The experts indexed by $\mathcal{V}_{i}$ of $\mathcal{G}_{i}$ share the most similar knowledge with each other and show much difference with experts indexed by $\mathcal{V}_{t}$ of $\mathcal{G}_{t}$, $t\neq i$. This can be formulated as the follows,
$\min\sum_{i=1}^{r}\left(\sum_{j,k\in\mathcal{V}_{i}}D_{jk}-\sum_{t\neq i}^{r}\sum_{j\in\mathcal{V}_{i},k\in\mathcal{V}_{t}}D_{jk}\right),$ (5)
[…]
where The objective function seeks to minimize the intra-group similarity (first term) and maximize the inter-group difference (second term) based on the pairwise similarity $D_{jk}$ based on expert $i$ and $j$.
Merging Similar Experts with Routers. To diversity the experts and preserve the different knowledge learned by different models clustered in the same group [37, 35], we merge the clusted experts with their routers on the weight space as follows,
$\hat{\theta}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}\theta_{\mathcal{V}_{n}(i)},~{}~{}\hat{W}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$ (6)
where $\mathcal{V}_{n}$ is the set of similar experts in the $n$-th cluster $\mathcal{G}_{n}$, and we have $\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}=1$. We update the MoE layer $F$ by respectively replacing all the experts grouped in $\mathcal{G}_{n}$ by a single FFN expert layer $f(\cdot;\hat{\theta}_{n})$ and corresponding routing weights $\hat{W}_{n}$, $n=1,2,...,N-r$.
3.4 Practical Implementation
On the discovery step. We propose two strategies for computing the disparity matrix $D$, including the data-centric strategy and model-centric strategy.
For data-centric strategy, we follow previous work [26] to use the pre-trained dataset C4, which can be a good surrogate to simulate the performance of the task-agnostic pruning. To improve the generalization ability at the discovery stage, we randomly mix the token embedding for data augmentation and then use the augmented data to calculate the expert output as the expert representation.
For model-centric strategy, we propose two ways to prune models with only expert weights, which encode the dataset information during the training process and can be a good agency for expert representation. The first way is to leverage the vectorized weights directly. The other way is to leverage the local linearity of neural networks to compute the surrogate weight matrix.
Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$. For input $x$, the output is computed as $f(x;\theta)=\theta_{2}(\sigma(\theta_{1}x)\cdot\theta_{3}x)$. Then, the vectorized weight is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$. Compared to vectorized weights, surrogate weights are more flexible with model size, showing more stable performance in pruning models (see section A.2).
On the merging step. We consider three approaches in practice deciding $\alpha$ to merge similar grouped experts. The first is to only maintain the experts with the maximum visiting frequency and drop all the others. The second is to set $\alpha_{i}=\frac{1}{|\mathcal{V}_{n}|}$, which uniformly assembles all the experts grouped. The last one is learning $\alpha$ to merge the grouped experts by minimizing the following loss function,
($\mathcal{L}(\{\alpha_{n}\}_{n=1}^{|\mathcal{V}_{n}|})=\|y-F(x;\hat{\Theta},\hat{W},K)\|,$ (7)
[…]
where the ground-truth is the output of original MoE layer $y=F(x;\Theta,W,K)$, and we jointly optimize $\lambda$ and $\alpha$ for different merging groups. Among these three strategies, we empirically find that the uniform souping strategy offers more stable performance and greater efficiency in model pruning (see section A.2 ). While the learning strategy can yield slightly better performance, it is more time-consuming due to the need for tuning the parameter $\alpha$.
4 Evaluation
Table 1: Results on pruning the Mixtral-8x7B from 8 experts to 6 and 4 experts in each MoE layer. The first and second columns respectively indicate the results of the pruned model with $6$ and $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
Router-guided [21] 51.8/24.8 60.5/26.5 46.9/24.7 60.5/25.0 82.6/39.9 32.0/11.6 70.4/50.9 57.8/29.1
Count-guided [11] 49.2/36.9 59.7/45.6 45.0/35.1 58.2/43.4 77.2/76.6 33.0/26.4 56.6/55.9 54.1/45.7
Enumerate [26] 52.4/43.5 66.4/52.7 49.0/40.4 63.7/43.5 84.0/80.8 32.6/28.8 71.1/66.4 59.9/50.8
Ours Model-centric 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
Data-centric 56.0/48.0 73.1/57.0 52.4/43.3 68.2/54.6 86.4/83.3 31.4/28.5 69.3/67.1 62.4/54.5
4.1 Experiment Setup
Studied Models. We take pruning three MoE-based architectures as an example, including the Mixtral, Deepseek, and Qwen. For the Mixtral architecture, the Mixtral-8x7B has $32$ sparse MoE-involved layers, in each there are $8$ experts. The Mixtral-8x22B is similar to Mixtral-8x7B but with $56$ sparse MoE layers. During the inference, each token will select $2$ experts in each MoE layer. The deepseek model has $28$ layers, and there are $64$ experts in each layer. Each token will pass $2$ shared experts and select $6$ experts during the inference. The Qwen model also has $28$ MoE layers with $64$ experts in each layer but will activate $8$ experts during the inference.
Pruning Methods. We take three advanced MoE pruning methods [11, 21, 26] as our baseline for comparison. Among them, router-guided merging [21] is initially designed for task-specific MoE pruning, where we set the target dataset as the samples from the pre-training dataset. For task-agnostic methods, we select the frequency-based pruning method Expert Trimming [26], also the count-guided strategy and loss-based pruning method [26], namely the Enumerate in our paper. Under the task-agnostic pruning setting, we disable all the fine-tuning stage of all methods for fair comparison.
For our method, we respectively report the best results of our data-centric and model-centric methods in pruning models. Following the previous setting [26], we use $128$ samples in C4 for computation in data-centric pruning methods, while the model-centric method doesn’t rely on the data. More detailed results on different implementations of discovery and merging steps (e.g., vectorized and surrogate weight strategy at the discovery step, max or learning strategy at the merging step) are provided in section A.2.
Table 2: Results on pruning the Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer. The first and second columns respectively indicate the results of the pruned model with $6$ and $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
Router-guided [21] 27.3/22.7 25.4/25.8 24.4/24.0 27.9/23.4 62.8/62.7 12.8/13.0 54.2/49.5 33.5/31.6
Count-guided [11] 58.0/45.7 74.9/57.7 54.1/42.0 70.2/45.7 81.5/74.4 35.2/27.0 69.3/57.4 63.3/50.0
Enumerate [26] 60.4/53.9 78.0/67.2 59.5/52.3 73.0/64.2 87.4/80.5 35.0/31.1 70.1/67.9 66.2/59.6
Ours Model-centric 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
Data-centric 62.3/57.8 78.5/69.7 60.2/51.3 73.4/64.2 87.6/83.1 35.8/33.2 71.1/68.1 67.0/61.1
Evaluation Datasets. The open-sourced Language Model Evaluation Harness library [10] is used to evaluate the performance. We select four zero-shot tasks, including MMLU [12], BoolQ [6], OpenBookQA [30], and RTE [1]. Among these tasks, MMLU consists of $57$ subtasks, where we present four groups, namely the humanities, social science, stem and other.
4.2 Pruning the Mixtral Architecture
We present the results of the pruning of the Mixtral architecture, including Mixtral-8x7B and Mixtral-8x22B. Both have $8$ experts in each MoE layer. We apply different methods to prune them from 8 experts to 6 and 4 experts in each layer, respectively. The results of the two models are shown in table 1 and table 2 respectively.
Results on Mixtral-8x7B. We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average. Compared with count-guided strategy [11] which just drops the experts less visited, router-guided strategy [21] has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge. Besides, compared with count-guided and enumerate strategies [26] which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers. We can also notice that the model-centric method surpasses all the other data-involved pruning baseline methods. This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning.
Results on Mixtral-8x22B. In this experiment, the model-centric method achieves the best result with only a minor performance drop of $2.8\%$ on average compared to the full model. Our proposed data-centric method achieves the second to the last performance, with a gap of $0.8\%$ performance gap compared with the runner-up method. It should be noted that the routing similarity-based pruning suffers model collapse under both two settings, showing a less robustness compared with other methods and metrics for similarity computation.
4.3 Pruning the Deepseek Model
We also evaluate the effectiveness of our proposed pruning method on compressing the Deepseek architecture. Specifically, we pruned the experts of Deepseek-MoE-16B, reducing the number of experts from 64 to 48 using different model pruning strategies. The results are illustrated in fig. 4.
Notably, even after pruning one-third of the experts in Deepseek-MoE-16B, our data-centric strategy maintains an impressive average performance of $50.9\%$, with only a $3.1\%$ reduction in performance compared to the full model. In the evaluation of the most challenging MMLU task, our model-centric strategy demonstrates superior performance in most cases, particularly when reducing the number of experts from $62$ to $52$. It consistently outperforms the runner-up baseline method, achieving a clear performance advantage of $2.6\%$.
Figure 4: Results on pruning the Deepseek-MoE-16B with different strategies.
4.4 Pruning the Qwen Model
Figure 5: Results on pruning the Qwen2-57B-14A with different strategies.
The Qwen architecture is also utilized in our experiments. We study the Qwen2-57B-14A to evaluate the performance of different model pruning strategies. We prune the experts from $64$ experts to $48$ experts in each layer and evaluate the performance on different tasks. The results are reported in fig. 5.
When pruning $\frac{1}{3}$ experts of the full model, the data-centric strategy delivers the best performance, outperforming both our model-centric method and the count-guided baseline method. On the MMLU task, our model-centric strategy achieves approximately $1.5\%$ better performance compared to the count-guided pruning strategy.
4.5 Empirical Analysis on Expert Hints
While expert pruning brings a significant speedup by reducing the parameters, what changes are introduced to different MoE layers? To answer this question, we analyze the change of expert hints on the calibration dataset C4.
Figure 6: Statistics of the visiting frequency for all experts in different MoE layers.
The statistic results are shown in fig. 6. We can see that in most MoE layers in Mixtral-8x7B, many pairs of experts have similar hints. In contrast, pruning differentiates the hints of experts. Compared to directly using hints as the pruning goal, using expert knowledge as the pruning criterion results in more significant changes in hints. Additionally, the merging operation on the gate aggregates the new expert (last column) more tokens, further increasing the hint differences.
Table 3: Evaluation results of different pruning setting with ranging the number of samples for expert similarity estimation. The augmentation is disabled in this experiment.
# of samples MMLU BoolQ OpenBookQA RTE Avg.
128 0.615 0.856 0.328 0.686 0.621
256 0.614 0.854 0.332 0.697 0.624
512 0.617 0.854 0.334 0.704 0.627
2,048 0.612 0.856 0.322 0.675 0.616
4.6 Ablation Study and Discussion
Comparison with the greedy search. To further demonstrate the superiority of our proposed expert pruning strategy, we employ a greedy search approach to identify the optimal candidates in each MoE layer for pruning and compare the results with those obtained using our method. Specifically, we enumerate all possible combinations of removing two experts in each MoE layer, evaluate the pruned models on the MMLU task, and record the best combination for each layer. We then merge these results to prune the model across all layers. The model pruned using the greedy search approach achieves a performance of $62.22\%$, while our data-centric strategy achieves $61.19\%$, showing a comparable result. We provide the full enumeration results in section A.1.
On the number of samples. We conduct experiments to study the effect of the used sample in data-centric pruning methods. As shown in table 3, when we increase the number of samples from $128$ to $512$, we can see a noticeable performance improvement with up to $0.6\%$. It should be noted that when we scale up the number of samples to $2,048$, the average performance dramatically drops to $61.6\%$.
We attribute this phenomenon to the use of mean operation in our expert similarity estimation. On the one hand, the mean operation keeps the feature dimension for efficient pairwise distance computation. On the other hand, it naturally downgrades the precision of function similarity computation by “round-off” error compared with other per sample-based distance metrics like CKA used in section 3.2.
On merging the routing policy. One relevant work to ours is [21], which identifies and merges similar experts based on routing policy, setting the routing policy of the merged expert to that of the most frequently visited expert. While we identify similar experts based on expert knowledge, such as weights and features, we merge not only the experts but also the routing policies. To demonstrate the importance of merging routers, we evaluate the performance of our pruning method without merging the routing policies. As shown in table 4, this results in performance degradation across all tasks and different numbers of experts, highlighting the crucial role of simultaneously merging the routers and experts.
Table 4: Evaluation results of our augmentation-based pruning method without merging the routers.
# of experts MMLU BoolQ OpenBookQA RTE Avg.
4 0.491 0.833 0.274 0.664 0.565
6 0.604 0.847 0.312 0.675 0.609
5 Conclusion and Future Work
In this paper, we address the practical problem of reducing the number of experts in MoE layers to accelerate the inference and fine-tuning process. We propose discovering similar experts at the feature level and then merging them in the weight space for MoE pruning while preserving as much original expert knowledge as possible. This approach allows the MoE layer to maintain diverse experts with different knowledge, thereby efficiently reducing redundancy.
However, current methods do not account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers. Our analysis reveals that some experts have low visitation frequencies, indicating a significant potential for further pruning.
6 Acknowledgments
We would like to thank Lifeng Li from Microsoft for setting up GPU machines.
Appendix A Evaluation on the expert similarity
Following the same setting in section 3.2, we conduct experiments on evaluating the expert similarity in all MoE layers of Mixtral-8x7B and Mixtral-8x22B. The results are respectively depicted in fig. A1 and fig. A2. We summarize the observation as follows,
  • Most MoE layers in the two Mixtral models contain significant expert redundancy.
  • The most redundant MoE layers are located in the first and last several layers, while the experts in the intermediate MoE layers learn more diverse features.
Figure A1: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under the linear kernel-based CKA criteria.
Figure A2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x22B under the linear kernel-based CKA criteria.
A.1 Enumeration on the expert pruning
We present the full greedy search result on Mixtral-8x7B in fig. A3. In detail, we first enumerate all the possible combinations of dropping $2$ experts layer by layer, and then evaluate the model on the MMLU task. For example, $66.19$ in the first row and third column in layer 0 indicates performance while dropping the first and third expert in layer 0 of the Mixtral-8x7B model.
Although dropping most of the combinations on different layers only leads to a minor performance drop, we can notice that it could cause the model to crash when the fourth expert in layer 1 is involved during the pruning process.
Figure A3: Enumeration on dropping two experts.
A.2 Results on different strategies for pruning
While we report the best performance using our data-centric and model-centric strategies to prune the MoE models, we detail more results on pruning the Mixtral-8x7B and Mixtral-8x22B.
For model centric, we show the results with vectorized weight and surrogate weight strategies to discover the similar experts. We uniformly merge the experts and their routers for model pruning. For data centric, after we discover similar experts, we respectively use the learning strategy to merge the experts with weights (Learn), or only maintain the expert with the maximum visiting frequency (Max).
From the results of our strategies, we can also notice two facts: 1) Both Vectorized $\theta$ and Surrogate $\theta$ strategies surpass all the other data-involved pruning baseline methods. This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning. 2) Compared with only using $\theta$, the involvement of $data$ in the pruning process can help better candidates for merging and pruning for improving the performance. Besides, though the learning strategy can bring a further improvement in the average performance, the computation cost is larger compared with using uniform coefficients for merging.
Table A1: Results on pruning the Mixtral-8x7B and Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq. 7; 3) Max: maintaining the expert in each discovered group with the maximum visiting frequency. The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
Model-centric Vectorized $\bf{\theta}$ 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
Surrogate $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
[…]
Max 56.4/47.6 71.9/58.6 52.0/42.9 66.9/55.7 85.1/82.8 35.2/28.8 70.4/66.8 62.6/53.7
Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
Model-centric Vectorized $\bf{\theta}$ 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
Surrogate $\bf{\theta}$ 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
[…]
Max $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7